# Understanding Self-Supervised Learning via Gaussian Mixture Models

### v2

@@ Abstract
Self-supervised learning attempts to learn representations from un-labeled data; it does so via a loss function that encourages the embedding of a point to be close to that of its augmentations. This simple idea performs remarkably well, yet it is not precisely theoretically understood why this is the case. In this paper we analyze self-supervised learning in a natural context: dimensionality reduction in Gaussian Mixture Models. Crucially, we define an augmentation of a data point as being another independent draw from the same underlying mixture component. We show that vanilla contrastive learning (specifically, the InfoNCE loss) is able to find the optimal lower-dimensional subspace even when the Gaussians are not isotropic – something that vanilla spectral techniques cannot do. We also prove a similar result for "non-contrastive" self-supervised learning (i.e., SimSiam loss). We further extend our analyses to multi-modal contrastive learning algorithms (e.g., CLIP). In this setting we show that contrastive learning learns the subset of fisher-optimal subspace, effectively filtering out all the noise from the learnt representations. Finally, we corroborate our theoretical finding through synthetic data experiments.
@@ 1 Introduction
Contrastive learning (CL) is now a gold-standard paradigm for learning representations with popular examples such as vision models like SimCLR Chen et al. (2020a) and MoCo Chen et al. (2020b) , text models like Dense Passage Retrieval (DPR) Karpukhin et al. (2020), and vision-language models like CLIP  Radford et al. (2021). Once learned, these representations both demonstrate commendable zero-shot performance, and could easily be adapted to achieve SoTA performance on various tasks like classification Chen et al. (2020a) and retrieval Izacard et al. (2021).
Essentially, contrastive learning is a self-supervised strategy that aims to learn consistent representations given a set of unlabeled dataset. The core idea is to relate similar (positive) data samples to each other while contrasting unrelated (negative) points to dissociate them. To formalize, consider a pair of points $({{\bm{x}}},\hat{{{\bm{x}}}})$, where ${{\bm{x}}}$ is some point from the dataset and $\hat{{{\bm{x}}}}$ is a “similar”, positive example. CL methods learn representations so that the embedding of the point ${{\bm{x}}}$ is close to that of its partner $\hat{{{\bm{x}}}}$, while being far away from the embedding of other points. Sometimes, these pairs are naturally present. For instance, in vision-language settings like CLIP, each training data comes as a pair of an image and the corresponding text caption. In other scenarios, they can be manually obtained via data augmentation that preserves the underlying semantics, e.g., we can define an image and its resized/cropped/rotated version as a pair.
It is fundamentally important to understand why this simple idea of relating/contrasting pairs works so well. In this paper, we aim to answer this question by a theoretical study of contrastive learning in a classic setting: dimensionality reduction in Gaussian Mixture Models (GMMs).
@@
To motivate our study for contrastive learning, let us take a particular representation learning perspective. Essentially, representation learning aims at finding a (non-convex) mapping between the high-dimensional input space of data, e.g., natural images of 1 million pixels, and a lower dimensional output domain of their representations, e.g., 1024-dimensional feature vectors. Similarly, contrastive learning aims at discovering robust mappings between the data and the representations by exploiting additional information in the form of augmentations. In our framework of GMMs, we consider linear mappings/projections (for analytical tractability) and characterize the optimality (more on this in Section 3) of the mapping functions learned by contrastive methods by analyzing the associated projection subspaces. An integral components of our work is a key, novel notion of augmentations in the context of GMMs.
@@
To highlight the success of contrastive methods in learning linear projectors, we conduct our study in comparison to standard spectral methods. i.e., based on singular value decomposition (SVD). We provide an example where spectral methods fall short of contrastive methods in finding anticipated projection subspaces (see Section 4.3). At this point, we would like to emphasize the main difference between contrastive methods and SVD-based methods; while contrastive learning methods leverage augmentations by design, the SVD-based methods cannot. Moreover, our endeavor is not to determine the better of the methods. In fact, we aim to show that when augmentations are present, contrastive learning can provably go beyond what would be possible otherwise. In that sense, our focus is on providing principled, theoretical insight into how contrastive learning works.
@@
Within our GMM framework, a key, novel proposal of ours is formalizing the “notion of augmentations”, i.e. point-pairs. We assume that for every point ${{\bm{x}}}$ drawn from the GMM, we are given an augmentation $\hat{{{\bm{x}}}}$, which is drawn from the same GMM based on a particular distribution (see Definition 5.1) that is biased towards the component that the original point ${{\bm{x}}}$ belongs to. To keep the setting realistic, we are oblivious to which GMM component either of them come from.
Given pairs of points $({{\bm{x}}},\hat{{{\bm{x}}}})$ of the type described above, we analyze contrastive learning objectives in two categories; single and multi-modal GMMs. In the single modal setting, the points in a pair are drawn from the same GMM distribution. Under this setting, we analyze InfoNCE and SimSiam losses, which are suited to leverage the additional information provided by the augmentations unlike spectral methods. In the multi-modal context, the pair $({{\bm{x}}},\hat{{{\bm{x}}}})$ will be generated by separate GMMs. A fundamental example is the CLIP model, where images are associated with text descriptions such that the samples are pairs of the form $({{\bm{x}}}_{V},{{\bm{x}}}_{T})$, where ${{\bm{x}}}_{V}$ (images) and ${{\bm{x}}}_{T}$ (texts) can have different dimensions. The objective now is to learn two different functions – two different linear projections – one for each modality, such that each modality is projected onto the same space. Contrastive learning in this context tries to make the final projection of ${{\bm{x}}}_{V}$ close to that of is pair ${{\bm{x}}}_{T}$, and far from random other $\tilde{{{\bm{x}}}}_{T}$’s. Particularly, the framework for the multi-modal GMM has a major difference; we do not need augmentations. We provide further details in Section 6.
@@
Contributions: Our objective is to theoretically understand the vital role augmentations in self-supervised representation learning; we do so in the context of GMMs. Precisely,
@@
    We formalize a mathematical notion of augmentation/point pairs – for both single and multi-modal GMMs, enabling us to concretely characterize contrastive learning methods in the context of GMMs.
@@
    We quantify a measure of optimality for the projections learned in the context of GMMs (see Section 4). We show that InfoNCE can find optimal projections for GMMs with general shared covariances, whereas spectral methods cannot do so beyond spherical GMMs.
@@
    For the multi-modal GMM setting, we show that the CLIP contrastive loss learns linear projections that are a subset of the (Fisher-)optimal subspaces for each modality, i.e., filtering out all noise directions.
2 Related Work
Contrastive Learning and InfoNCE.
@@ 2 Related Work
Contrastive learning is widely used in current state-of-the-art representation learning methods like SimCLR Chen et al. (2020a), MoCo Chen et al. (2020b) etc. Prior theoretical work analysing contrastive learning argue that the contrastive objective leads to learning of good "general purpose" representations that can be used for better downstream performance and sample complexity. Arora et al. (2019) gives the first theoretical analyses of the contrastive loss, showing provable guarantees on downstream task with reduced sample complexity. HaoChen et al. (2021) relaxes the augmentation assumptions made in Arora et al. (2019) by showing that contrastive learning is equivalent to spectral clustering on some appropriately defined data graph. Saunshi et al. (2022); HaoChen and Ma (2022) extend this line of work by arguing for the inclusion of the inductive biases of the neural network architecture into the theoretical analyses.
Self-supervised Learning.
@@
Apart from contrastive methods, BYOL Grill et al. (2020), SwAV Caron et al. (2020), SimSiam Chen and He (2021) are also popular methods for self-supervised representation learning. These methods are also based on Siamese architecture used in contrastive methods, but notably don’t rely on negatives to prevent a "collapsing" solution. Contrastive methods like SimCLR Chen et al. (2020a) and CLIP Radford et al. (2021) aim to learn representations so that embedding of a point is close to its partner in the pair while far away from embedding of a random other point. The non-contrastive self-supervised learning methods enforce just the former condition. These methods empirically work by breaking the symmetry between the Siamese towers Wang et al. (2022). Prediction head, momentum encoder, and stop gradient operation are some the of the common ways to achieve this asymmetry. Prior work  Tian et al. (2021); Jing et al. (2021); Wen and Li (2022) aim to understand the training dynamics of and the role of each of the operation in preventing collapse. We instead focus on the fixed point analyses of the loss functions.
Dimensionality Reduction for GMMs.
@@
Spectral Methods are popularly used for dimensionality reduction in gaussian mixture models. Spectral clustering algorithms Sanjeev and Kannan (2001); Vempala and Wang (2004); Achlioptas and McSherry (2005); Kannan et al. (2009); Brubaker and Vempala (2008) pair these spectral methods with standard clustering algorithms (e.g. K-Means) in the low dimensional space. Since the error bounds of these methods vary as square root of the ambient dimension, spectral clustering following the two step process leads to lower clustering error than clustering in the original space itself.
Our work does not focus on state-of-the-art methods for dimensionality reduction for Gaussian mixtures, but instead leverages the problem of dimensionality reduction to analyze self-supervised learning. Specifically, we aim to analyze the role of augmentations/positives in this setup.
@@ 3 Problem Setup
We consider data distributions following a Gaussian mixture model (GMM). Intuitively, each mixture component in GMM represents a class in data.
Definition 3.1 (SharedGMM).
A Shared Covariance Gaussian Mixture Model parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$ is defined as the probability distribution $F=\sum_{k\in[K]}w_{k}\mathcal{N}({\bm{\mu}}_{k},{\bm{\Sigma}})$ where $\{{\bm{\mu}}_{k}\}_{k\in[K]}$ are the means and ${\bm{\Sigma}}\in\mathbb{S}_{++}^{d}$ is the positive definite covariance matrix shared by the components. $w_{k}$ are the positive mixture weights summing to one.
Task Formulation.
@@
We consider the task of learning a “good” representation function $f(.)$ for points ${{\bm{x}}}$ drawn from a data distribution. Informally, such a function should map the data onto a lower-dimensional subspace while preserving the class information. For tractability, we restrict $f$ to the class of linear mappings; ${{\bm{x}}}\mapsto{{\bm{A}}}^{T}{{\bm{x}}}$, where ${{\bm{A}}}\in\mathbb{R}^{d\times r}$. In this setup, we will show that self-supervised methods can learn the optimal mapping for this task. The standard approach for learning such a mapping is using spectral algorithms Kannan et al. (2009) (See Sec 4.2 for details). We compare and contrast our results for self-supervised methods to well-known results for spectral methods.
Spectral vs Self-supervised Methods.
Our results show that self-supervised methods are “better” than spectral methods. We highlight two keys points which contribute to the better performance of self-supervised methods. First, spectral methods learn the linear mappings assuming as input a dataset of points, i.e., ${{\bm{x}}}\sim\mathcal{D}$. In comparison, self-supervised methods operate on input data given as pairs i.e. $({{\bm{x}}},\hat{{{\bm{x}}}})\sim\mathcal{D}_{\mathrm{pair}}$ where ${{\bm{x}}}$ and $\hat{{{\bm{x}}}}$ are considered to be similar. Second, their optimization problems differ in their target objectives. While spectral objectives find the best-fit subspace explaining maximum data variance, self-supervised objectives are based on the principle of bringing mappings of similar point closer (in conjunction with a loss-specific regularizer).
${\mathcal{L}_{\mathrm{spectral}}({{\bm{A}}};\mathcal{D})=\underbrace{-\ ||{{\bm{A}}}^{T}{{\bm{x}}}||^{2}}_{\mathrm{(negative)\ explained\ variance/best\ fit}}}$
${\mathcal{L}_{\mathrm{selfsup}}({{\bm{A}}};\mathcal{D}_{\mathrm{pair}})=-\underbrace{({{\bm{A}}}^{T}{{\bm{x}}})^{T}{{\bm{A}}}^{T}\hat{{{\bm{x}}}}}_{\mathrm{attractive\ term}}+\underbrace{\mathcal{R}({{\bm{x}}})}_{\mathrm{regularization}}}$
We elaborate on these differences in more detail in Sec 5.
3.1 Linear Dimensionality Reduction (LDR)
We direct the reader’s attention to the similarity between our task and linear dimensionality reduction (LDR) for Gaussian mixtures. Framing our task as linear dimensionality reduction allows us to establish a quantitative measure of comparison between mappings learned by spectral (${{\bm{A}}}_{\mathrm{spectral}}$) and self-supervised methods (${{\bm{A}}}_{\mathrm{selfsup}}$).
@@ 3.1 Linear Dimensionality Reduction (LDR)
For a SharedGMM, we denote it’s "intra-component" variance (i.e. variance within a component) by ${\bm{\Sigma}}$ and it’s "inter-component" variance (i.e. the distance between components) by $\sum_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}^{T}_{k}$ (following  Fukunaga (2013)). A GMM having a low intra-component variance and a high inter-component variance is said to be well-separated. LDR aims to map high dimensional GMMs into well-separated low dimensional GMMs. Fisher discriminant Fukunaga (2013) evaluates the performance of such mappings.
@@
Suppose a SharedGMM (Def 3.1) parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$. Then, the fisher discriminant $J({{\bm{A}}})$ for a mapping ${{\bm{A}}}$ (mapping the SharedGMM) is defined as:
$J({{\bm{A}}})=\mathrm{Tr}\Bigg{(}\bigg{[}{{\bm{A}}}^{T}{\bm{\Sigma}}{{\bm{A}}}\bigg{]}^{-1}\bigg{[}{{\bm{A}}}^{T}\big{(}\sum_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T}\big{)}{{\bm{A}}}\bigg{]}\Bigg{)}$
Note that ${{\bm{A}}}^{T}{\bm{\Sigma}}{{\bm{A}}}$ and ${{\bm{A}}}^{T}(\sum_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T}){{\bm{A}}}$ denote the intra-component and the inter-component variances respectively, after the projection operation via ${{\bm{A}}}$. Thus, the Fisher discriminant measures the ratio of inter-component to intra-component variances after projection.
@@
A “favorable” mapping ${{\bm{A}}}$ will have a high value of the Fisher discriminant $J({{\bm{A}}})$. Hence, the Fisher discriminant $J$ serves as the quantitative metric for evaluating ${{\bm{A}}}_{\mathrm{spectral}}$ and ${{\bm{A}}}_{\mathrm{selfsup}}$.
4 Fisher Subspace and Spectral Methods
SVD Projection Contrastive Learning
Data Single point ${{\bm{x}}}\sim F$ Pair of points $({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}$
Objective Maximize data variance Min. intra-component variance
Max. inter-component variance
Fisher-optimality Spherical GMMs Shared Covariance GMMs
Table 1: We highlight the differences between classic SVD-based linear dimensionality reduction paradigm and our self-supervised (contrastive) learning-inspired problem setting.
4.1 Fisher Subspace
It is known that for any target dimension $r$, the mapping having the top $r$ eigenvectors of ${\bm{\Sigma}}^{-1}(\sum_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T})$ as its columns, maximizes the Fisher discriminant. We can also deduce that for any two mappings ${{\bm{A}}}_{1}$ and ${{\bm{A}}}_{2}$ such that $\mathrm{Col}({{\bm{A}}}_{1})\subseteq\mathrm{Col}({{\bm{A}}}_{2})$, we have $J({{\bm{A}}}_{1})\leq J({{\bm{A}}}_{2})$, i.e., $J(\cdot)$ is monotonic in the subspace of the matrix. Based on this observation, we will have the following definition, which is fundamental in quantifying the optimality of projections learned by different methods.

### v3

# Understanding Contrastive Learning via Gaussian Mixture Models

@@ Abstract
Contrastive learning involves learning representations via a loss function that encourages each (unlabeled) sample to be far from other samples, but close to its own augmentation. In this paper, we aim to understand why this simple idea performs remarkably well, by theoretically analyzing it for a simple, natural problem setting: dimensionality reduction in Gaussian Mixture Models (GMMs). Note that the standard GMM setup lacks the concept of augmentations. We study an intuitive extension: we define the pair of data sample and its augmentation as a coupled random draw from the GMM such that the marginal over the "noisy" augmentation is biased towards the component of the data sample. For this setup, we show that vanilla contrastive loss, e.g., InfoNCE, is able to find the optimal lower-dimensional subspace even when the Gaussian components are non-isotropic. In particular, we show that InfoNCE can match the performance of a fully supervised algorithm, e.g., LDA, (where each data point is labeled with the mixture component it comes from) even when the augmentations are "noisy". We further extend our setup to the multi-modal case, and develop a GMM-like setting to study the contrastive CLIP loss. We corroborate our theory with experiments on CIFAR100; representations learned by InfoNCE loss match the performance of LDA on clustering metrics.
@@ 1 Introduction
Contrastive learning (CL) is now a gold-standard paradigm for learning representations with popular examples such as vision models like SimCLR [6] and MoCo [8] , text models like Dense Passage Retrieval (DPR) [18], and vision-language models like CLIP  [21]. Once learned, these representations both demonstrate commendable zero-shot performance, and could easily be adapted to achieve SoTA performance on various tasks like classification [6] and retrieval [13].
[…]
@@
It is fundamentally important to understand why this simple idea of relating/contrasting pairs works so well. In this paper, we aim to answer this question with a theoretical study of contrastive learning by casting is as dimensionality reduction for Gaussian Mixture Models (GMMs). We motivate our study through the lens of representation learning which aims at finding a mapping between the high-dimensional input space, e.g., images of 1 million pixels, and a lower dimensional output domain of representations, e.g., 1024-dimensional feature vectors. Similarly, contrastive learning aims at discovering robust mappings between the data and the representations by exploiting additional information in the form of augmentations. In our framework of GMMs, we consider linear mappings/projections (for analytical tractability) and characterize the optimality of the mapping functions learned by contrastive methods by analyzing the associated projection subspaces.
@@
To highlight the success of contrastive methods in learning linear projectors, we conduct our study in comparison to fully supervised methods and standard spectral methods, i.e., based on singular value decomposition (SVD). We provide an example where we provide a three-way comparison in finding anticipated projection subspaces (see Section˜3.1). Note that SVD-based methods are not designed to leverage augmentations. Therefore, we use SVD-based methods merely as a reference point to argue that when (noisy) augmentations are present, contrastive learning can provably go beyond what would otherwise be possible. However, a more relevant and interesting comparison is with respect to supervised methods, i.e., LDA in this case. In that sense, we provide principled, theoretical insight into how contrastive learning works and how it performs with respect to a measure of optimality.
@@
Within our GMM framework, a key proposal of ours is formalizing the “notion of augmentations”, i.e. point-pairs, going beyond the idealized definition considered in [3]. In the context of GMMs, a natural choice of augmentation would be drawing a random sample from the same component as the original point belongs to. Instead, we assume that for every point $\bm{x}$}}, its augmentation $\hat{{{\bm{x}}}}$ is a random draw from the GMM following a distribution (see Definition˜2.2) that is biased towards the component of the original point $\bm{x}$}}. To keep the setting realistic, we are oblivious to which GMM component either of them comes from.
[…]
@@
Contributions: Our objective is to theoretically understand the vital role of augmentations in self-supervised representation learning; we do so in the context of GMMs. To summarize,
@@
    We formalize a generalized notion of noisy augmentation/point pairs, enabling us to concretely characterize contrastive learning methods in the context of GMMs.
@@
    We quantify a measure of optimality for the projections learned in the context of GMMs. We show that InfoNCE can find optimal projections for GMMs with general shared covariances.
@@
    For the multi-modal setting, we show that the CLIP loss learns linear projections that are a subset of the (Fisher-)optimal subspaces for each modality, filtering out noise directions.
1.1 Related work
@@
Contrastive learning and InfoNCE. Contrastive learning is at the heart of state-of-the-art representation learning methods like SimCLR [6], MoCo [8] etc. Prior theoretical work on contrastive learning argue that the contrastive objective leads to learning of "general purpose" representations that facilitates better downstream performance and sample complexity. [2] gives the first theoretical analyses of the contrastive loss, showing provable guarantees on downstream task with reduced sample complexity. [12] relaxes the augmentation assumptions made in [2] by showing that contrastive learning is equivalent to spectral clustering on some appropriately defined data graph. [23, 11] extend this line of work by arguing for the inclusion of the inductive biases of the neural network architecture into the theoretical analyses. [16, 28] extend the setup of [12] to show that the eigenfunctions of the positive pair graph (as defined in [12]) are the basis of the set of view-invariant functions. They further show that these representations are equivalent to Kernel PCA with “positive-pair kernel”. [20] show that under the condition that the augmentation sets within clusters are close and intertwined (relative to the complexity of the learning function class), the InfoNCE loss is able to learn consistent representations.
@@
Self-supervised learning. BYOL [10], SwAV [5], SimSiam [7] are some of the most popular methods for self-supervised representation learning. They are also based on Siamese architecture used in contrastive methods, but notably, they do not make use of negatives to prevent “collapsing” solutions. Contrastive methods like SimCLR [6] and CLIP [21] aim to learn representations so that embedding of a point is close to its partner while far away from that of any other point. On the contrary, non-contrastive self-supervised learning methods only enforce the former condition and do not explicitly deal with the separation between the embeddings of samples. These methods are empirically motivated by the notion of breaking the symmetry between the encoders in Siamese networks [26]. [24, 15, 27] propose various techniques for breaking this symmetry and study their training dynamics to understand their role in preventing collapse. Compared to the relevant work on self-supervised learning, we instead focus on the fixed point analysis of the loss functions.
@@
Dimensionality reduction for GMMs. Spectral methods are popularly used for dimensionality reduction in GMMs by analyzing the principal directions of the data covariance. Spectral clustering algorithms [22, 25, 1, 17, 4] essentially combine spectral analysis with standard clustering algorithms (e.g. K-Means) in the low dimensional space. Since their error bounds grow as square root of the ambient dimension, the two-step approach leads to lower clustering error than in the original space.
@@ 2 Problem setup
We consider the task of learning a “good” representation function for points $\bm{x}$}} drawn from a data distribution. Informally, such a function should map the data onto a lower-dimensional subspace while preserving the class information. For tractability, we restrict $f$ to the class of linear mappings; ${{\bm{x}}}\mapsto{{\bm{A}}}^{T}{{\bm{x}}}$, where ${{\bm{A}}}\in\mathbb{R}^{d\times r}$. In this setup, we will show that self-supervised methods can learn the optimal mapping for this task. We compare and contrast our results for self-supervised methods to well-known results for supervised methods (e.g., LDA) and spectral methods.
As the basis of our study, we consider data distributions following a Gaussian mixture model (GMM). Intuitively, each mixture component in GMM represents a class in data and we assume that components have a shared covariance.
Definition 2.1.
A Shared Covariance Gaussian Mixture Model (SharedGMM) parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$ is defined as the probability distribution $F=\sum_{k\in[K]}w_{k}\mathcal{N}({\bm{\mu}}_{k},{\bm{\Sigma}})$ where $\sum_{k}w_{k}=1$, ${\bm{\mu}}_{k}$ are the means and ${\bm{\Sigma}}\in\mathbb{S}_{++}^{d}$ is the covariance matrix shared by the components.
To complete our data model, we formalize the definition of augmentation. Given a point $\bm{x}$}}, its augmentation $\hat{{\bm{x}}}$ is an independent sample from the mixture distribution with a bias towards the underlying component of $\bm{x}$}}. The bias implies that if $\bm{x}$}} is sampled from a component $z$, then its augmentation, on average, is more likely to be sampled from the same component $z$.
Definition 2.2.
For a SharedGMM $F$ parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$, we define its Augmentation-enabled Distribution (AeD) $\hat{F}$, parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}},\delta\}_{k\in[K]}$ as
$\hat{F}=\delta\sum\nolimits_{k}w_{k}\big(\mathcal{N}({\bm{\mu}}_{k},{\bm{\Sigma}})\times\mathcal{N}({\bm{\mu}}_{k},{\bm{\Sigma}})\big)+(1-\delta)\big(\sum\nolimits_{k}w_{k}\mathcal{N}({\bm{\mu}}_{k},{\bm{\Sigma}})\times\sum\nolimits_{k^{\prime}}w_{k^{\prime}}\mathcal{N}({\bm{\mu}}_{k^{\prime}},{\bm{\Sigma}})\big)$
To elucidate, augmentation-enabled distribution (AeD) returns a pair of points $({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}$ in a two-step process. We first flip a coin with bias $\delta$. Based on the coin flip, we either sample twice from the same component or sample from possibly different components (chosen with probability $w_{k}$). The bias of the coin $\delta$ controls the correlation between $\bm{x}$}} and $\hat{{{\bm{x}}}}$: $\delta=0$ means $\bm{x}$}} and $\hat{{{\bm{x}}}}$ are independent samples, while $\delta=1$ means $\bm{x}$}} and $\hat{{{\bm{x}}}}$ are independent draws from the same component. The definition ensures that the marginal of $\bm{x}$}} and $\hat{{{\bm{x}}}}$ are equal to $F$ for any $\delta$.
Spectral versus self-supervised methods. Our results corroborate the known advantages of self-supervised methods over spectral methods. Clearly, spectral methods learn linear mappings directly on the data samples ${{\bm{x}}}\sim\mathcal{D}$, whereas, self-supervised methods operate on the augmented pairs $({{\bm{x}}},\hat{{{\bm{x}}}})\sim\mathcal{D}_{\mathrm{pair}}$. Consequently, their optimization problems differ in their target objectives:
[…]
While spectral objectives find the best-fit subspace explaining maximum data variance, self-supervised objectives are based on the principle of bringing embeddings of similar point closer in conjunction with a loss-specific regularizer. The regularizer typically induces separation between distinct points.
@@
Linear dimensionality reduction (LDR). Framing our task as LDR allows us to establish a measure of comparison between mappings learned by spectral (${{\bm{A}}}_{\mathrm{spectral}}$) and self-supervised methods (${{\bm{A}}}_{\mathrm{selfsup}}$). For a SharedGMM, we denote it’s "intra-component" variance (i.e. variance within a component) by $\bm{\Sigma}$} and "inter-component" variance (i.e. the separation between components) by $\sum_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}^{T}_{k}$ [9]. A GMM is said to be well-separated if it has low intra-component variance and high inter-component variance. LDR aims to map a GMM to a low dimensional subspace where it is well-separated. We evaluate the performance of these mappings by Fisher discriminant [9].
@@
Suppose a SharedGMM (Def 2.1) parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$. Then, the fisher discriminant $J({{\bm{A}}})$ for a mapping $\bm{A}$}} is defined as:
$J({{\bm{A}}})=\mathrm{Tr}\left(\left[{{\bm{A}}}^{T}{\bm{\Sigma}}{{\bm{A}}}\right]^{-1}\left[{{\bm{A}}}^{T}\left(\sum\nolimits_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T}\right){{\bm{A}}}\right]\right)$
@@
Note that ${{\bm{A}}}^{T}{\bm{\Sigma}}{{\bm{A}}}$ and ${{\bm{A}}}^{T}(\sum_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T}){{\bm{A}}}$ denote the intra-component and the inter-component variances, respectively, after projection via $\bm{A}$}}. Thus, the Fisher discriminant measures the ratio of inter-component to intra-component variances post-projection. A “favorable” mapping $\bm{A}$}} will have a high value of the Fisher discriminant $J({{\bm{A}}})$, hence, the Fisher discriminant $J$ serves as the quantitative metric for evaluating the subspaces learned by different frameworks.
Connection to linear discriminant analysis (LDA). One can argue that it is not surprising for self-supervised methods to have better performance than classical spectral methods. The more interesting comparison would be against a supervised method, which would help uncover the capabilities of contrastive methods. Therefore, we focus our attention on Linear Discriminant Analysis (LDA) [9]. LDA is a supervised LDR method that leverages the class information to learn the subspace where the data distribution conditioned on the class label (i.e. underlying component index) is maximally separated (see Appendix˜H). We highlight that assuming AeD (Def. 2.2) is strictly weaker than assuming a labeled dataset. We show that this relaxed condition is enough for learning the Fisher subspace and gives similar performance to supervised LDR methods.
3 Fisher subspace
For any target dimension $r$, the linear map that has the top $r$ eigenvectors of ${\bm{\Sigma}}^{-1}(\sum_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T})$ as its columns maximizes the Fisher discriminant. Similarly, for any mappings ${{\bm{A}}}_{1}$ and ${{\bm{A}}}_{2}$, if $\mathrm{Col}({{\bm{A}}}_{1})\subseteq\mathrm{Col}({{\bm{A}}}_{2})$, then we have $J({{\bm{A}}}_{1})\leq J({{\bm{A}}}_{2})$. In other words, $J(\cdot)$ is monotonic in the subspace of the matrices. Based on this, we will have the following definition, which is necessary in measuring the optimality of projections learned by different methods.